OpenStudy (abbycross167):

can someone please help me answer this question? I will fan and give a medal!! Question~ write the following inequalities in y = mx + b form. –3x –4y > 12

OpenStudy (notshh):

it just wants you to get y by itself. so add 3x to both sides and then divide both sides by -4

OpenStudy (abbycross167):

@notshh so to add 3x to both sides it will look like this? (-3)–3x –4y > 12(-3) ?

OpenStudy (notshh):

adding 3x (not -3x) to both sides would give you -4y >12 + 3x then from there, divide -4 from both sides. what do you get?

OpenStudy (anonymous):

change the inequality also,because when we multiply or divide by a negative quantity sign of inequality changes.

OpenStudy (notshh):

^ ye

OpenStudy (abbycross167):

@notshh I am very sorry for not replying sooner!! I'm so sorry! If you don't mind helping me still I would really like your help.. Please... ~~~ Can you check my steps when I divide by -4? Divide both sides by -4 (-4) -4y>12+3x(-4) (-4)/-4=1 3/(-4)= 0.75 = 1y>12+0.75x

OpenStudy (mathmale):

–3x –4y > 12 is to be solved for 4y, and then for y alone. You may find it easier to multiply this whole inequality by (-1), remembering to reverse the direction of the > symbol. You should get 3x+4y<12. Subtract 3x from both sides. Show your work. Divide both sides of that result by 4. This is the inequality you wanted, in the form y<mx+b.

OpenStudy (notshh):

I'm not really following your work. D: But the final answer you got looks as if you divided by a positive 4 instead of -4. so from here –4y > 3x + 12 after dividing by -4, the inequality sign (>) will be flipped to <, and the numbers to the right of the sign will change to negative.

OpenStudy (abbycross167):

oh ok, could you try to show me what it looks like when dividing? Like the steps you do? Because I'm a little confused on what it should look like...

OpenStudy (notshh):

just divide each term by -4 -4y ÷ -4 = ? 3x ÷ -4 = ? 12 ÷ -4 = ?

OpenStudy (abbycross167):

-4y ÷ -4 = 1 3x ÷ -4 = -0.75 12 ÷ -4 = -3 Then what?

OpenStudy (mathmale):

You may choose which approach you prefer: 1) multiply the given inequality by (-1) and also reverse the direction of the > symbol, or 2) don't bother to change the given inequality by (-1). Since you and notshh have started out by following the latter approach, Add 3x to both sides of the equation to move the term -3x to the right side: –3x –4y > 12 +3x = 3x -------------- -4y>12+3x Next, divide each of these 3 terms by -4 to isolate y, and change the direction of the > sign. -4y > 12 + 3x ----- --- --- -4 -4 -4 Change the direction of the inequality sign now. Next, simplify. Abby, your turn!

OpenStudy (mathmale):

\[\frac{ -4y }{ -4 }=?\]

OpenStudy (mathmale):

\[\frac{ 12 }{ -4 }=?\]

OpenStudy (mathmale):

\[\frac{ 3x }{-4 }=?\]

OpenStudy (notshh):

now that you know all the values, you can write the inequality. don't forget to flip the inequality sign since you divided by a negative.

OpenStudy (mathstudent55):

@mathmale –3x –4y > 12 Multiplying the inequality above by -1 on both sides gives 3x + 4y < -12 (It's -12, not 12)

OpenStudy (abbycross167):

@mathmale so all I have to do is simplify the fractions ?

OpenStudy (mathmale):

That's right. But watch the signs carefully.

OpenStudy (praavi):

Step 1: Multiply by '-' (negative symbols) on both sides of the equation Why? It makes the equation have positive signs as ('-'x'-' is '+') Looks like after multiplying by negative sign: 3x+4y<-12 Observe the inequality symbol ( changed from > to <) Why does it change? Assume: 5>4 If I multiply both sides by -1 and keep the symbol as same, it is an error -5>-4 As we all know, -5 is lesser than -4. Hence the above is incorrect, so we flip the inequality Got it? Step 2: Subtract 3x on both sides We are concentrating on the form y=mx+b So, we see that y is singled out on LHS So, lets keep 'y' on one side of the '=' symbol. Oops, but we've got 4y. So, keep '4y' on LHS. Subtract 3x on both sides of the equation as 3x is added on both sides. Looks like after subtracting: 4y<-12-3x Step 3: Now, our aim is to get only 'y'. So, we divide by 4 on both sides as it is multiplied with 'y'. Now, the answer looks like after dividing by 4 on both sides: y<(-12-3x)/4 I hope you get it, I 'll respond if you have any more doubts regarding this if I have time.

OpenStudy (abbycross167):

-4/-4=1 12/-4= -3 3/-4= -3/4 @mathmale

OpenStudy (praavi):

That's right

OpenStudy (mathmale):

-4/-4=1 is actually -4y/(-4) = ? (Can't drop the variable.) Your other two divisions turned out fine.

OpenStudy (abbycross167):

so it would be 1y?

OpenStudy (mathmale):

Now put all three terms into the appropriate places in one inequality. Either 1y, or simply y.

OpenStudy (mathmale):

Abby: Note the appropriate cautionary message I received from mathstudent55: "–3x –4y > 12 Multiplying the inequality above by -1 on both sides gives 3x + 4y < -12 (It's -12, not 12)" He's right. Take this into account as y ou write your final answer (inequality).

OpenStudy (mathmale):

All you'll need to do (following his point) is to change one sign in your inequality.

OpenStudy (abbycross167):

oh ok, Thank you all so much!!! I understand it now! Thank y'all again!

OpenStudy (mathmale):

My pleasure. Thanks, mathstudent55, praavi and notshh.